# Activation-Informed Merging of Large Language Models

### v2

@@ 3.4 Adaptable Relaxation Scheme
As discussed, we introduce an adaptable relaxation scheme based on the activations of the base model, which we wish not to stray away from significantly. To make the scheme adaptable to any merging algorithm in the weight space, we formulate our relaxation scheme in terms of the changes in the weights. Given a task-agnostic representative calibration corpus . We can pass this corpus through the model and accumulate the activations from each token in the dataset. This will yield the average magnitude of activations for each channel in all layers of the model. As previously noted, averaging the inputs over the calibration set yields a vector for each linear layer. Constructing a diagonal matrix from the absolute values of this vector gives . By normalizing using its maximum value, we obtain a diagonal matrix . This matrix serves to modulate weight updates based on input saliency, allowing for consistent control of merging through a universal hyperparameter . Next, we define the action of any given merging method by the changes that it applies to the model weights with respect to each of the fine-tuned models being merged (i.e., ). Specifically, we denote the weight update contributed by a fine-tuned model (with parameters ) to the model parameters by (e.g, for weight averaging). Now, we propose an adaptive relaxation scheme that adjusts the final model. For simplicity in notation let refer to weights of a linear layer (i.e. an matrix), then the AIM relaxation scheme can be written as:
@@
Note that in general is not a single matrix rather a mapping of activations for all model parameters obtained as we described prior. In our experiments, we apply this relaxation scheme to several different merging methods and explore how the hyperparameter affects the merged model’s behavior, and present these results in Section 5.
@@
An alternative way of choosing the importance scores for changing the model weights—which has been used, e.g., in continual learning Farajtabar et al. (2020); Min et al. (2022), out-of-distribution detection Sharma et al. (2021), and meta-learning Almecija et al. (2022)—is to use the partial derivatives of the (base) model with respect to the parameters, i.e., sensitivities, which correspond to a scaled version of the activations. This is in contrast with using the activations directly, common in model compression. Formally, let denote the model with parameters , and be one of the weights; then determines how sensitive the output is to perturbing the weight . We develop a sensitivity-based formulation of AIM by incorporating sensitivity scores calculated from the gradients and replacing activations with sensitivity scores. We consider to be the logits of the model with parameter and to be the entropy function, similar to Farajtabar et al. (2020)’s suggested approach for classification problems. Now let be the magnitude of gradients for all parameters of the pretrained model for samples in a calibration corpus . Then the sensitivity-based formulation can be written as follows:
In other words, the sensitivity-based formulation of the problem suggests that regularization of weights through relaxation should be done based on model gradients. However, we note that computing gradients for a calibration corpus can be computationally expensive and will require significantly more memory and compute resources than storing activations, which is akin to performing inference on the model. To verify that using activations retains the same performance and fidelity as this formulation, we conduct our experiments for both AIM and the sensitivity-based formulation. We present the results of the continual learning view in Appendix C, and we see that when comparing results of AIM (Table 1) and the sensitivity-based formulation (Table 3) the performance boost of both approaches is very similar, while AIM being computationally more efficient.
@@ 4.1 Selection of FT Expert LLMs
To understand how AIM reacts with different merging methods, we conduct experiments with merging different experts fine-tuned from the same base model. The set of experts we use is the same set of experts used by the two latest LLM merging algorithms in the literature, namely DARE Yu et al. (2024a) and WIDEN Yu et al. (2024b), which use the same three experts fine-tuned from Llama-2-13b Touvron et al. (2023). These experts include the WizardLM-13B Xu et al. (2024) model fine-tuned for instruction following, WizardMath-13B Luo et al. (2025) fine-tuned for superior mathematical reasoning, and llama-2-13b-code-alpaca which serves as the code expert Chaudhary (2023). See Appendix A for more details.
@@ 4.2 Merging Methods Implementations
In our experiments, we implement the latest merging methods in the literature for LLM merging. These include newly developed DARE and WIDEN Yu et al. (2024a, b) methods as well as some of the long-established approaches of TIES merging Yadav et al. (2023), and task arithmetic Ilharco et al. (2023). For all merging methods except WIDEN, we use the comprehensive MergeKit implementations developed by Goddard et al. (2024), and for WIDEN we use the publicly available implementation provided by the authors of the paper.
We note that in many of the merging algorithms, many hyper-parameters can be adjusted. In these cases, we use the author-recommended values where available and the default parameters recommended by Goddard et al. (2024). Note that it is possible to perform a grid search on these hyper-parameters to find optimal values for each benchmark, however, this would essentially be over-fitting on benchmarks and does not provide any value to our analysis of the proposed complementary relaxation scheme which applies adjustments to the merged models. For reproducibility, all of our checkpoints and code to reproduce the results will be made publicly available.
@@ 4.3 Benchmarks Used For Evaluations
Given that the expert models we use in our experiments involve fine-tuning on instruction following, mathematical reasoning, and code generation we use several common benchmarks for each of these tasks. Specifically, we measure model performance on language understanding with the MMLU Hendrycks et al. (2021a) benchmark, instruction following with IFEval Zhou et al. (2023) benchmark, code generation with HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) benchmarks, and mathematical reasoning with the MATH Hendrycks et al. (2021b) and GSM8K Cobbe et al. (2021) benchmarks. For all benchmarks, we use the latest versions and up-to-date implementations developed by Gao et al. (2024) except for mathematical reasoning, for which we use the chain of thought prompting used by Luo et al. (2025) to replicate the results of the original model as closely as possible. The code we use for these benchmark results will also be made publicly available for reproducibility.
@@ 4.4 Measuring Performance From an Optimization Perspective
Where is the number of dimensions/benchmarks. Since hypervolume is computed only over Pareto-optimal models, we have , ensuring that . This metric provides an aggregated measure of merging effectiveness, capturing trade-offs across multiple benchmarks rather than focusing on isolated improvements thus providing a full picture of merging performance. In our experiments, we track HV Gain along with the 6 aforementioned benchmarks.
5 Experiments & Results
@@ 5.1 AIM Applied to Various Merging Approaches
To demonstrate the effectiveness of AIM, we conduct experiments on 5 different merging methods under 4 different scenarios. As mentioned before we use 3 different FT LLM experts in our experiments. As such we merge models using each merging method for all 4 possible permutations of these expert LLMs. Then we apply AIM to all merged models and measure the performance of each model in all 6 benchmarks. We also report the HV gain for each merged model compared to the population of the base model and the models being merged (in cases with 2 models the population will only include the models used for merging). These results are presented in Table 1. For this experiment, we used , which we found to be the best balance of performance among the various merging methods we use. This choice was informed by our analysis in Section 5.2. In Table 1 we have highlighted the gain/loss of performance for each benchmark due to AIM and we can see that in the vast majority of cases, AIM causes a significant performance boost, with an Average Percentage-Point Change of 13% (ignoring the Inf value) and more than 40% HV Gain in 20% cases, further highlighted by the fact that the top performers on each benchmark, as well as the largest hypervolume gain, are all in models merged with AIM. We observe HumanEval (10 out of 20) and MBPP (17 out of 20) often see large boosts with AIM, especially when merging Instruction Tuned models with others. Some merges also reveal small drops in GSM8K or IFEval even when other benchmarks improve, reflecting the inherent trade-offs in merging specialized models. Overall, a clear majority (80%) of merges exhibit improved HV Gain under AIM, reinforcing that the method often enhances multi-task performance overall. We can further visualize this increase in hypervolume by looking at how AIM pushes the Pareto frontier. Figure 2 shows how applying AIM to existing merging methods extends the Pareto optimal frontier, which we also quantitatively measured using HV gain. These results showcase the efficacy of the proposed method across a variety of merging methods and reinforce the hypothesis that the activation space encompasses useful insight for merging.
@@ 5.2 Ablation study
To understand the effects of changing in AIM, we conduct an ablation study on the case of merging all three expert LLMs. For this study, we apply AIM with and run the benchmarks on each merged model with each value of . For brevity, we do not report all benchmark results for each value here, instead, we track the hypervolume gain (The full set of results are presented in Appendix B). Specifically, to better visualize the effect of , we measure the relative change in HV Gain compared to no AIM (i.e. ). We present these results in Figure 3. In most merging methods, we see that decreasing to even still benefits the model performance while, in some cases, still increases the performance. However, in TIES merging particularly, we see that decreasing beyond seems to degrade performance, and setting to the most extreme case of does see some degradation in WIDEN as well. Given this, it seems that in these experiments, a value of balances the performance gains in methods responding well to AIM and the potential degradation of methods that benefit less from AIM. However, given this observation that in some cases pushing to still yields benefits, there may be some value in exploring non-linear scaling of activation magnitudes and non-linear relaxation schemes that could further boost performance in some cases.
@@ 6 Conclusion and Outlook
Through extensive empirical evaluations across multiple merging methods and benchmark tasks, we demonstrated that AIM consistently improves performance, often yielding superior results in comparison to the original merging methods it was applied to. These results empirically confirm our hypothesis on the importance of the activation space. Notably, AIM boosted merged model performance by up to 40% in some cases, underscoring the crucial role and the potential of activation information in merging methods. Our findings highlight the necessity of incorporating activation-informed strategies when merging multiple fine-tuned models.
Moving forward, our findings open up several promising directions for future research. First, our results indicate that even aggressively preserving salient weights of the pre-trained model is effective across many merging scenarios. This highlights the promise for more advanced activation-informed strategies and non-linear relaxation methods to enhance performance potentially further. Beyond the pre-trained activations explored in this work, there is room to improve existing merging methods by leveraging the broader activation space of the models being merged. So far, AIM has only considered the activations of the pre-trained model, while the activations of the expert LLMs remain unexplored. Future research should focus on developing methods that also encompass information from the expert model activations. Additionally, in future works, more theoretically grounded approaches for incorporating the activation space of LLMs in merging should be developed and tested. These integrations will hold great value in improving the quality and performance of merging methods in an increasingly competitive and ever more efficient landscape of LLMs, which could benefit from smaller and more efficient yet more powerful models.
Overall, AIM serves as a robust and adaptable augmentation to existing LLM merging techniques, offering a principled way to incorporate activation information for more effective model fusion. By prioritizing the activation-aware perspective, we take a step towards more stable, efficient, and generalizable merged models that better leverage the strengths of multiple fine-tuned experts.
@@ Appendix A Reproducibility Details
A Note On Weights: The weights we use in our experiments may not be exactly identical to the weights used in the experiments by Yu et al. (2024a) and Yu et al. (2024b), since the referenced weights for WizardMath-13B are no longer available publicly, instead we use a publicly available copy of the model.
Code & Data:
Aside from the checkpoint, we provide our code and the link to the publicly available calibration data we use in our work. Our code is publicly available at https://anonymous.4open.science/r/ActivationInformedMerging and the calibration data can be found at https://huggingface.co/datasets/mit-han-lab/pile-val-backup.

### v3

@@ 3.4 Adaptable Relaxation Scheme
As discussed, we introduce an adaptable relaxation scheme based on the activations of the base model, which we wish not to stray away from significantly. To make the scheme adaptable to any merging algorithm in the weight space, we formulate our relaxation scheme in terms of the changes in the weights. Given a task-agnostic representative calibration corpus . We can pass this corpus through the model and accumulate the activations from each token in the dataset. This will yield the average magnitude of activations for each channel in all layers of the model. As previously noted, averaging the inputs over the calibration set yields a vector for each linear layer. Constructing a diagonal matrix from the absolute values of this vector gives . By normalizing using its maximum value, we obtain a diagonal matrix . This matrix serves to modulate weight updates based on input saliency, allowing for consistent control of merging through a universal hyperparameter . Next, we define the action of any given merging method by the changes that it applies to the model weights with respect to each of the fine-tuned models being merged (i.e., ). Specifically, we denote the weight update contributed by a fine-tuned model (with parameters ) to the model parameters by (e.g., for weight averaging). Now, we propose an adaptive relaxation scheme that adjusts the final model. For simplicity of notation, let refer to weights of a linear layer (i.e., an matrix), then the AIM relaxation scheme can be written as:
@@
Note that in general is not a single matrix, rather a mapping of activations for all model parameters obtained as we described prior. In our experiments, we apply this relaxation scheme to several different merging methods and explore how the hyperparameter affects the merged model’s behavior, and present these results in Section 5.
@@
An alternative way of choosing the importance scores for changing the model weights—which has been used, e.g., in continual learning [farajtabar2020OGD, min2022one], out-of-distribution detection [sharma2021sketching], and meta-learning [almecija2022uncertainty]—is to use the partial derivatives of the (base) model with respect to the parameters, i.e., sensitivities, which correspond to a scaled version of the activations. This is in contrast with using the activations directly, common in model compression. Formally, let denote the model with parameters , and be one of the weights; then determines how sensitive the output is to perturbing the weight . We develop a sensitivity-based formulation of AIM by incorporating sensitivity scores calculated from the gradients and replacing activations with sensitivity scores. We consider to be the logits of the model with parameter and to be the entropy function, similar to farajtabar2020OGD’s suggested approach for classification problems. Now let be the magnitude of the gradient for all parameters of the pretrained model for samples in a calibration corpus . Then the sensitivity-based formulation can be written as follows:
In other words, the sensitivity-based formulation of the problem suggests that regularization of weights through relaxation should be done based on model gradients. However, we note that computing gradients for a calibration corpus can be computationally expensive and will require significantly more memory and compute resources than storing activations, which is akin to performing inference on the model. To verify that using activations retains the same performance and fidelity as this formulation, we conduct our experiments for both AIM and the sensitivity-based formulation. We present the results of the continual learning view in Appendix C, and we see that when comparing results of AIM (Table 1) and the sensitivity-based formulation (Table 4), the performance boost of both approaches is very similar, with AIM being computationally more efficient.
@@ 4.1 Selection of FT Expert LLMs
To understand how AIM reacts with different merging methods, we conduct experiments with merging different experts fine-tuned from the same base model. The set of experts we use is the same set of experts used by the two latest LLM merging algorithms in the literature, namely DARE [DARE] and WIDEN [widen], which use the same three experts fine-tuned from Llama-2-13b [llama2]. These experts include the WizardLM-13B [xu2024wizardlm] model fine-tuned for instruction following, WizardMath-13B [wizardmath] fine-tuned for superior mathematical reasoning, and llama-2-13b-code-alpaca, which serves as the code expert [codealpaca]. See Appendix A for more details.
@@ 4.2 Merging Methods Implementations
In our experiments, we implement the latest merging methods in the literature for LLM merging. These include newly developed DARE and WIDEN [DARE, widen] methods as well as some of the long-established approaches of TIES merging [yadav2023tiesmergingresolvinginterferencemerging], and task arithmetic [TaskArithmetic]. For all merging methods except WIDEN, we use the comprehensive MergeKit implementations developed by goddard-etal-2024-arcees, and for WIDEN, we use the publicly available implementation provided by the authors of the paper.
We note that in many of the merging algorithms, many hyperparameters can be adjusted. In these cases, we use the author-recommended values where available and the default parameters recommended by goddard-etal-2024-arcees. Note that it is possible to perform a grid search on these hyperparameters to find optimal values for each benchmark; however, this would essentially be over-fitting on benchmarks and does not provide any value to our analysis of the proposed complementary relaxation scheme, which applies adjustments to the merged models. For reproducibility, all of our checkpoints and code to reproduce the results will be made publicly available.
@@ 4.3 Benchmarks Used For Evaluations
Given that the expert models we use in our experiments involve fine-tuning on instruction following, mathematical reasoning, and code generation, we use several common benchmarks for each of these tasks. Specifically, we measure model performance on language understanding with the MMLU [hendrycks2020measuring] benchmark, instruction following with IFEval [zhou2023instructionfollowingevaluationlargelanguage] benchmark, code generation with HumanEval [chen2021codex] and MBPP [austin2021program] benchmarks, and mathematical reasoning with the MATH [hendrycks2021measuringmathematicalproblemsolving] and GSM8K [gsm8k] benchmarks. For all benchmarks, we use the latest versions and up-to-date implementations developed by lmharness except for mathematical reasoning, for which we use the chain of thought prompting used by wizardmath to replicate the results of the original model as closely as possible. The code we use for these benchmark results will also be made publicly available for reproducibility.
@@ 4.4 Measuring Performance From an Optimization Perspective
where is the number of dimensions/benchmarks. Since hypervolume is computed only over Pareto-optimal models, we have , ensuring that . This metric provides an aggregated measure of merging effectiveness, capturing trade-offs across multiple benchmarks rather than focusing on isolated improvements, thus providing a full picture of merging performance. In our experiments, we track HV Gain along with the 6 aforementioned benchmarks.
5 Experiments and Results
@@ 5.1 AIM Applied to Various Merging Approaches
To demonstrate the effectiveness of AIM, we conduct experiments on 5 different merging methods under 4 different scenarios. As mentioned before, we use 3 different FT LLM experts in our experiments. As such, we merge models using each merging method for all 4 possible permutations of these expert LLMs. Then we apply AIM to all merged models and measure the performance of each model in all 6 benchmarks. We also report the HV gain for each merged model compared to the population of the base model and the models being merged (in cases with 2 models, the population will only include the models used for merging). These results are presented in Table 1. For this experiment, we used , which we found to be the best balance of performance among the various merging methods we use. This choice was informed by our analysis in Section 5.2. In Table 1, we have highlighted the gain/loss of performance for each benchmark due to AIM and we can see that in the vast majority of cases, AIM causes a significant performance boost, with an Average Change of 13% (ignoring the Inf value) and more than 40% HV Gain in 20% cases, further highlighted by the fact that the top performers on each benchmark, as well as the largest hypervolume gain, are all in models merged with AIM. We observe HumanEval (10 out of 20) and MBPP (17 out of 20) often see large boosts with AIM, especially when merging Instruction Tuned models with others. Some merges also reveal small drops in GSM8K or IFEval even when other benchmarks improve, reflecting the inherent trade-offs in merging specialized models. Overall, a clear majority (80%) of merges exhibit improved HV Gain under AIM, reinforcing that the method often enhances multi-task performance overall. We can further visualize this increase in hypervolume by looking at how AIM pushes the Pareto frontier. Figure 2 shows how applying AIM to existing merging methods extends the Pareto optimal frontier, which we also quantitatively measured using HV gain. These results showcase the efficacy of the proposed method across a variety of merging methods and reinforce the hypothesis that the activation space encompasses useful insight for merging.
To validate AIM’s generalizability beyond the Llama-2 family, we conducted a new experiment on a different architecture and modality: the Qwen 2.5-VL-7B vision-language model. We merged two distinct experts, Video-R1 model for video reasoning [feng2025videoR1] and CAD-Coder model for image-to-code-based CAD geometry generation [doris2025cadcoder], with their instruct base model.
We evaluated this merge on 6 benchmarks: IFEval and MMLU for instruct base model capabilities, Video MMMU [hu2025videommmu] and VSI-Bench [yang2025vsi] for video reasoning, and two CAD-Coder benchmarks for code generation from rendered and real images [doris2025cadcoder]. As shown in Table 2, AIM consistently improves the underlying merge methods across these diverse tasks and improves the Hypervolume Gain under all four merging methods. This demonstrates that AIM’s benefits are not architecture-specific and generalize effectively to multimodal models.
For merging, therefore, we have two experts, one for CAD and one for video reasoning, with a base model that is instruction-tuned. Therefore, to assess performance, we perform the IF-Eval and MMLU benchmarks (the base model knowledge and instruction following) as well as expert benchmarks for video reasoning, Video MMMU [hu2025videommmu] and VSI-Bench [yang2025vsi], and CAD generation benchmarks of CAD-Coder benchmark (on rendered CAD images) [doris2025cadcoder] and CAD-Coder Real benchmark (on real 3D printed images) [doris2025cadcoder]. As shown in Table 2 below, AIM consistently improves the performance of the underlying merging method across diverse tasks while also showing that the overall hypervolume gain is raised in all merging scenarios, demonstrating AIM’s benefits are not limited to the Llama-2 architecture, and AIM does provide a path towards higher quality merging with little computational and data overhead.
| Method | AIM | IFEval | MMLU | Video VSI | Video MMMU | CAD test100 | CAD R400 | HV Gain |
| Base Models | ||||||||
| Instruct | - | 0.68 | 0.67 | 0.21 | 0.47 | 0.04 | 0.05 | - |
| Video-R1 | - | 0.64 | 0.61 | 0.38 | 0.49 | 0.00 | 0.03 | - |
| CADCoder | - | 0.27 | 0.66 | 0.00 | 0.00 | 0.61 | 0.32 | - |
| Merged Models | ||||||||
| TIES | No | 0.57 | 0.68 | 0.23 | 0.41 | 0.61 | 0.32 | 0.44 |
| Yes | 0.58 (+1.75%) | 0.68 | 0.23 | 0.45 (+9.76%) | 0.65 (+6.56%) | 0.32 | 0.45 (+2.26%) | |
| DARE | No | 0.54 | 0.67 | 0.25 | 0.42 | 0.54 | 0.32 | 0.43 |
| Yes | 0.56 (+3.70%) | 0.67 | 0.26 (+4.00%) | 0.45 (+7.14%) | 0.55 (+1.85%) | 0.32 | 0.44 (+3.47%) | |
| DARE Task Arithmetic | No | 0.25 | 0.64 | 0.25 | 0.43 | 0.54 | 0.33 | 0.38 |
| Yes | 0.56 (+124.00%) | 0.67 (+4.69%) | 0.26 (+4.00%) | 0.45 (+4.65%) | 0.54 | 0.34 (+3.03%) | 0.45 (+17.42%) | |
| WIDEN | No | 0.35 | 0.66 | 0.08 | 0.26 | 0.61 | 0.33 | 0.32 |
| Yes | 0.35 | 0.66 | 0.14 (+75.00%) | 0.32 (+23.08%) | 0.64 (+4.92%) | 0.33 | 0.36 (+13.24%) | |
@@ 5.2 Ablation study
To understand the effects of changing in AIM, we conduct an ablation study on the case of merging all three expert LLMs. For this study, we apply AIM with and run the benchmarks on each merged model with each value of . For brevity, we do not report all benchmark results for each value here; instead, we track the hypervolume gain (The full set of results are presented in Appendix B). Specifically, to better visualize the effect of , we measure the relative change in HV Gain compared to no AIM (i.e., ). We present these results in Figure 3. In most merging methods, we see that decreasing to even benefits the model performance. However, in TIES merging particularly, we see that decreasing beyond seems to degrade performance, and setting to the most extreme case of does see some degradation in WIDEN as well. Given this, it seems that in these experiments, a value of balances the performance gains in methods responding well to AIM and the potential degradation of methods that benefit less from AIM. However, given this observation that in some cases pushing to still yields benefits, there may be some value in exploring non-linear scaling of activation magnitudes and non-linear relaxation schemes that could further boost performance in some cases.
To empirically validate the sensitivity to the calibration set, we conducted an ablation study on the calibration set size. We applied AIM to the DARE TIES merge of all three Llama-2 experts from Table 1, varying the number of calibration blocks from 1 to 256. As shown in Figure 4, the results demonstrate that AIM is highly robust. The Hypervolume Gain (HV Gain) increases substantially with just one block and stabilizes by 8 blocks (sourced from the Pile corpus gao2020pile800gbdatasetdiverse). This confirms that the data overhead for AIM is minimal and the method is highly practical.
@@ 6 Conclusion and Outlook
Through extensive empirical evaluations across multiple merging methods and benchmark tasks, and model architectures (including the Llama-2 and Qwen-VL families), we demonstrated that AIM consistently improves performance, often yielding superior results in comparison to the original merging methods it was applied to. These results empirically confirm our hypothesis on the importance of the activation space. Notably, AIM boosted merged model performance by up to 40% in some cases, underscoring the crucial role and the potential of activation information in merging methods. Furthermore, our ablation study confirmed AIM’s robustness, showing significant gains even with minimal calibration data (as few as 8 blocks), highlighting the method’s practical applicability. Our findings strongly highlight the necessity and benefit of incorporating activation-informed strategies when merging multiple fine-tuned models.
Moving forward, our findings open up several promising directions for future research. First, our results indicate that even aggressively preserving salient weights of the pre-trained model is effective across many merging scenarios. This highlights the promise for more advanced activation-informed strategies and non-linear relaxation methods to potentially further enhance performance. Beyond the pre-trained activations explored in this work, there is room to improve existing merging methods by leveraging the broader activation space of the models being merged. So far, AIM has only considered the activations of the pre-trained model, while the activations of the expert LLMs remain unexplored. Future research should focus on developing methods that also encompass information from the expert model activations. Additionally, in future works, more theoretically grounded approaches for incorporating the activation space of LLMs in merging should be developed and tested. These integrations will hold great value in improving the quality and performance of merging methods in an increasingly competitive and ever more efficient landscape of LLMs, which could benefit from smaller and more efficient yet more powerful models.
Overall, AIM serves as a robust and adaptable augmentation to existing LLM merging techniques, offering a principled way to incorporate activation information for more effective model fusion. By prioritizing the activation-aware perspective, we take a step towards more stable, efficient, and generalizable merged models, demonstrated across different architectures and modalities, that better leverage the strengths of multiple fine-tuned experts.
Acknowledgments and Disclosure of Funding
This work was supported in part by the MIT-IBM Watson AI Lab.
@@ Appendix A Reproducibility Details
A Note On Weights: The weights we use in our experiments may not be exactly identical to the weights used in the experiments by DARE and widen, since the referenced weights for WizardMath-13B are no longer available publicly, instead we use a publicly available copy of the model.
Code and Data:
Aside from the checkpoint, we provide our code and the link to the publicly available calibration data we use in our work. Our code is publicly available at https://github.com/ahnobari/ActivationInformedMerging and the calibration data can be found at https://huggingface.co/datasets/mit-han-lab/pile-val-backup.